Three dice are rolled at once what is the probability of obtaining same face faces
1/36
1/2
1/6
1/72

Answer by greenestamps(13200)



Imagine rolling the three dice one at a time.

The first die can be any number: P = 1
The second die has to be the same as the first: P = 1/6
The third die also has to be the same as the first: P = 1/6

P(all three the same) = (1)(1/6)(1/6) = 1/36